OpenStudy (anonymous):

Can some one help me Solve this please. A weight of Water is 62 1/2lb per cubic foot. What is the weight of 6 1/8 cubic feet of water. How much does 6 1/8 cubic feet of water weight?

OpenStudy (anonymous):

Convert both of your fractions to proper ones and multiply\[62\frac{ 1 }{ 2 }=\frac{ 125 }{ 2}\] \[6\frac{ 1 }{ 8 }=\frac{ 49 }{ 8 }\]You multiply because you are converting units. It's like this: when you have x pounds per cubic foot and you multiply by y cubic feet, the cubic feet cancel and you are left with xy pounds.

OpenStudy (anonymous):

Does this sound right?